# Probing modified Hawking evaporation with gravitational waves from the primordial black hole dominated universe

### v1

@@ I Introduction
However, it has been recently argued that some of the conclusions based on black hole evaporation can be evaded. Ref. [12] examined “quasi-extremal” PBHs, for which the evaporation efficiency is reduced, and showed that a longer PBH lifetime (and associated smaller luminosity) makes them evade bounds and potentially remain a feasible dark matter candidate. Refs. [13] and [14] have explored PBHs in the framework of large extra dimensions [15] and demonstrated that this allows for new mass windows for light PBHs as dark matter candidates.
@@
For 4-dimensional black holes, Ref. [16] argued that Hawking’s semi-classical calculations fail at the latest when the black hole has shed about half of its initial mass. Their main argument is that Hawking’s result completely disregards the backreaction of the emission on the black hole itself. Refs. [17, 16] then suggest that this effect should no longer be negligible when the energy of the emitted quanta becomes comparable to that of the black hole itself. The proposed backreaction leads to a “memory burden” [17], which effectively suppresses the evaporation rate by inverse powers of the black hole entropy $S^{-n}$. Since for $M>M_{\rm pl}$ the entropy of a black hole is “large”, the memory burden can significantly extend the black hole’s lifetime. Interestingly, Refs. [18, 19] recently noted that, in the presence of the memory burden, PBHs with masses below $10^{10}\rm g$ could still account for the whole DM.
@@
In this work, we focus on even lighter PBHs which may have dominated the energy density of the universe in its very early stages and subsequently led to reheating via the emission of thermal radiation through Hawking evaporation [20, 21, 22, 23, 24]. Crucially, as shown in Refs. [25, 26, 27, 28, 29, 30, 31] (see also Refs. [32, 33] for reviews), a PBH dominated stage in the very universe could lead to an observable GW background. This signature would result from perturbations induced by PBH shot noise at formation, while second-order induced GWs associated with the PBH formation mechanism would fall at frequencies above the LIGO/Virgo/Kagra band [34] in the so-called ultra-high frequency range [35, 36]. Since the GW production in this scenario depends on the dynamics before and after PBH evaporation, it may represent a unique opportunity to test modifications of Hawking evaporation. In this work, we will show that GWs induced right after evaporation by PBH number density fluctuations have the potential to probe the memory burden model. Also, since the main effect of the memory burden is to extend the duration of the PBH domination, but still with a fast final burst of PBHs but does not alter much the final instants of evaporation, the GW signal is enhanced with respect to the standard case.
@@ II Memory burden and the PBH dominated universe
where $M_{\rm mb}=qM_{\rm f}$ parametrizes the moment when the modification to the semi-classical approximation sets in, with $M_{\rm f}$ being the PBH mass at formation and $q$ a free parameter. We introduced ${\cal M}$ to distinguish two different possibilities discussed in Ref. [19]:
@@
where $g_{H}(T_{\rm PBH})$ are the spin-weighted degrees of freedom heavier than the PBH temperature $T_{\rm PBH}=M_{\rm pl}^{2}/M$. In what follows, we only consider ultra-light PBHs with $M\ll 10^{11}\,{\rm g}$ and, therefore, we use the high temperature SM relativistic degrees of freedom $g_{H}(T_{\rm PBH})\approx 108$, unless stated otherwise. The semi-classical (sc) evaporation is recovered by taking the limit $n\to 0$. Note that we use the subscript “f” to denote evaluation at the time of formation.
@@
We note that our results are also applicable to the scenarios considered in Ref. [12]. In particular, we can underline a few relevant examples. We introduce $\varepsilon$ which parametrizes, in their notation, the departure from the standard Hawking evaporation of a Schwarzschild black hole leading to a temperature $T\sim\frac{[\varepsilon/(1+\varepsilon)^{2}]}{M_{\rm PBH}}$. In the case of a nearly extremal Reissner-Nordstrom charged black hole, $\varepsilon=\sqrt{1-Q^{2}/M^{2}}\ll 1$, where $Q$ represents the PBH charge. To consider these scenarios, it is sufficient to assume $q=1$ while identifying the factor $\alpha^{-1}q^{3+2n}S_{\rm f}^{n}$ with $\frac{(1+\varepsilon)^{6}}{2^{6}\varepsilon^{4}}$. Another notable example is provided by higher dimensional black holes [13, 14]. An initial BH of size $r_{s}$ much larger than the size of extra dimensions $R$ evaporates following the 4d decay equation until its horizon becomes $r_{s}\simeq R$. Afterward, it would continue its decay at a rate that is slower than 4D due to the effect of the extra dimension. We can derive the corresponding phenomenological quantities by relating Eq. (4) to Eq. (A22) in [39].
@@ II.1 Extended PBH reheating
Let us now turn to the PBH reheating scenario (see e.g. Refs. [25, 27, 28] for more details on the standard case). We assume that PBHs form from the collapse of large primordial fluctuations after inflation in the radiation dominated universe. For analytical simplicity, we further assume a monochromatic PBH mass spectrum (for a broad mass spectrum in the PBH reheating scenario, see Ref. [30]) with the PBH mass given by
@@
$\Gamma=-\frac{{\rm d}\log M}{{\rm d}t}=\begin{cases}\frac{1}{3(t_{\rm eva}^{\rm sc}-t)}\qquad{\rm for}\quad t\leq t_{\rm mb}\\ \frac{1}{3\alpha(t_{\rm eva}-t)}\qquad{\rm for}\quad t>t_{\rm mb}\end{cases}~{}.$ (12)
@@
In general, we find two possibilities depending on whether the early PBH-radiation equality (which we denote as “eeq” and which is defined by $\rho_{\rm PBH}(t_{\rm eeq})=\rho_{\rm rad}(t_{\rm eeq})$) happens before or after the memory burden time $t_{\rm mb}$ (7). In the first case, that is $t_{\rm mb}<t_{\rm eeq}$, the initial fraction of the PBHs is too low to reach matter radiation equality before the memory burden time. In fact, this case roughly corresponds to the standard scenario where PBHs evaporate before dominating the universe. Now, due to the memory burden, PBHs have a second chance to dominate the universe. In the second case, where $t_{\rm mb}>t_{\rm eeq}$, the early PBH-radiation equality occurs before the memory burden time. Then, as seen in Fig. 1, there is a short period in the PBH dominated stage when the PBH energy density decays, but PBHs still dominate the universe. We find that, in that case, $\Omega_{\rm PBH}(t_{\rm mb})\approx 0.59$ for $q=1/2$. It should be noted that the precise value of $\Omega_{\rm PBH}(t_{\rm mb})$ depends on $q$. For example, for $q\lesssim 0.41$ we find that $\Omega_{\rm PBH}(t_{\rm mb})<1/2$ and radiation temporarily dominates the universe. Depending on the parameteres of the modified evaporation, we find that for small enough values of $q$ PBHs do not come to dominate again after $t_{\rm mb}$, while for other choices there is a second stage of PBH domination. We do not consider these possibilities here and focus on the regime $10.41<q<$ 1. In particular, we will take $q=\frac{1}{2}$ for all plots, unless stated otherwise. Also, note that the PBH remnant scenario would correspond to $q\sim M_{\rm pl}/M_{\rm f}$ (and $n\to\infty$ so that Hawking evaporation stops completely).
@@
$M_{\rm f}<M^{\rm md}_{\rm f,max}=5\times 10^{8}\,{\rm g}\left(\frac{e^{12n}q^{n+\frac{3}{2}}7^{\frac{2n}{3}}10^{8n}}{\alpha}\right)^{-\frac{2}{2n+3}}\approx 5\times 10^{8}\,{\rm g}\left(\frac{e^{31.7n}q^{n+\frac{3}{2}}}{\alpha}\right)^{-\frac{2}{2n+3}}~{}.$ (19)
@@
Lastly, we can express the BBN bound in terms of an upper bound on $n$ by enforcing the Hubble parameter at formation to be smaller than the aforementioned bound $H_{\rm f}<2.5\times 10^{-5}\,M_{\rm pl}$ [41], which provides a lower bound for the mass which is roughly given by $M_{\rm f,max}^{\rm md}\gtrsim 1\,{\rm g}$. Thus, we find that
$n<\frac{59-3\log q}{2(12+\log q)}\,.$ (20)
@@ III Memory burden and the induced GW signal from PBH density fluctuations
We proceed to estimate the GW spectrum induced by PBH number density fluctuations. We will closely follow Refs. [28, 32], which builds upon the work of Refs. [25, 46, 47, 26, 27, 32], adapting the relevant calculations to the current case. The main effect of the modification to Hawking evaporation given by Eqs. (4) is an extension of the PBH dominated era, proportional to the PBH entropy. Thus, one naively expects that the amplitude of the GW spectrum is enhanced and that the peak moves to lower frequencies with respect to the standard case.
@@
$\frac{k_{\rm uv}}{k_{\rm eva}}\approx 4989\,\alpha^{-1/3}q^{1/3}e^{8n}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2}{3}(1+n)}\,,$ (22)
In the case where $t_{\rm mb}<t_{\rm eeq}$, one should add a factor $q^{-1/3}$ to the right hand side of Eq. (22).
@@
${f_{\rm uv}}\approx 3.7\times 10^{6}\,{\rm Hz}\,\alpha^{1/6}q^{1/3}e^{-4n}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{-\frac{5}{6}-\frac{n}{3}}\left(\frac{g_{*}(T_{\rm eva})}{106.75}\right)^{1/4}\left(\frac{g_{*s}(T_{\rm eva})}{106.75}\right)^{-1/3}\,,$ (24)
@@
${\cal P}_{\Phi}(k)\Big{|}_{\rm eva}={\cal A}_{\Phi}\left(\frac{k}{k_{\rm uv}}\right)^{m}\Theta(k_{uv}-k)\,,$ (25)
@@
We will use $\Omega_{\rm GW,\rm res}^{\rm peak}$ from Eq. (III) as a good approximation for the peak amplitude of the GW spectrum. It should be noted that in the case when $\beta<\beta_{*}$, there are additional suppression factors. First, there is a $q^{8}$ coming from $k_{\rm eeq}$ in Eq. (III) and one $q^{-1/3(7-4/(3\alpha))}$ from $k_{\rm uv}/k_{\rm eva}$, leading to a total suppression of $q^{\frac{17}{3}+\frac{4}{9\alpha}}$. [^1] The ratio $k_{\rm uv}/k_{\rm eva}$ gets an additional $q^{-1/3}$ dependence because the time of equality $t_{\rm eeq}$ is greater than in the case $\beta>\beta_{*}$. There is one factor $q$ coming from $a_{\rm f}/a_{\rm eeq}$ and a factor $q^{-4/3}$ from $a_{\rm eeq}/a_{\rm eva}$, see Eq. (15). Taking $q\sim 1/2$ and $\alpha\sim 1$ gives an additional suppression of $0.015$, which just implies a factor $2$ larger value for $\beta$. For simplicity, we ignore this effect. The GW spectrum evaluated today is then given by
$\Omega_{\rm GW,0}h^{2}$ $\approx 1.62\times 10^{-5}\left(\frac{\Omega_{r,0}h^{2}}{4.18\times 10^{-5}}\right)\left(\frac{g_{*}(T_{\rm rh})}{106.75}\right)\left(\frac{g_{*s}(T_{\rm rh})}{106.75}\right)^{-4/3}\Omega_{\rm GW,eva}(f)\,.$ (32)
@@
Figure 5: The induced gravitational wave spectrum (32) for PBH mass ${M_{\rm f}=10^{3}$\mathrm{g}$}$, $\beta=3\times 10^{-8}$ and $n=\frac{1}{2}$ in the scenario mb1 ($\alpha=1$) (solid red line), as well as the spectrum for scenario mb2 ($\alpha=1+2n/3$) with a PBH mass ${M_{\rm f}=2\times 10^{3}$\mathrm{g}$}$, $\beta=2\times 10^{-12}$ and $n=1$ (dashed red line). As one can see, for this choice of parameters, the signals enter the observational windows for several of the next-generation GW detectors. We also show the power-law integrated sensitivity curves as in Ref. [50] (but also see Ref. [51]) for LIGO A+, Voyager, Einstein Telescope (ET) (triangular configuration with 15km arms [52]), Cosmic Explorer (CE) and DECIGO experiments (for the sensitivity curves see Refs. [53, 54, 55, 51, 56, 57]). In purple, we show the upper bounds from the search for an isotropic GW background from the LVK collaboration [58]. The current (integrated) constraint from BBN [59, 60, 61] is shown with a dashed blue line.
@@
$\Omega_{\rm GW,\rm res}^{\rm peak}\approx 1.2\times 10^{15}\,\beta^{16/3}q^{4}\left(\frac{\sqrt{2}\alpha}{\sqrt{3\alpha-1}}\right)^{-\frac{4}{3\alpha}}\frac{e^{\frac{8}{\alpha}\left[{(7\alpha-\frac{4}{3})n+\frac{17}{12}(\alpha-1)}\right]}}{\left({\alpha}/{q}\right)^{\frac{7}{3}-\frac{4}{9\alpha}}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2(21\alpha-4)(1+n)}{9\alpha}}\,.$ (34)
@@
$\beta<1.2\times 10^{-4}\,q^{-3/4}\frac{\alpha^{\frac{1}{6\alpha}+\frac{7}{16}}q^{\frac{1}{12\alpha}-\frac{7}{16}}}{(3\alpha-1)^{\frac{1}{8\alpha}}}e^{\frac{-84\alpha n+16n+17}{8\alpha}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{-\frac{(21\alpha-4)(1+n)}{24\alpha}}\,.$ (35)

### v2

@@ I Introduction
However, it has been recently argued that some of the conclusions based on black hole evaporation can be evaded. Ref. [12] (see also [13]) examined “quasi-extremal” PBHs, for which the evaporation efficiency is reduced, and showed that a longer PBH lifetime (and associated smaller luminosity) makes them evade bounds and potentially remain a feasible dark matter candidate. Refs. [14] and [15, 16] have explored PBHs in the framework of large extra dimensions [17] and demonstrated that this allows for new mass windows for light PBHs as dark matter candidates.
@@
For 4-dimensional black holes, Ref. [18] argued that the latest instant Hawking’s semi-classical calculations can break down is when the black hole has shed about half of its initial mass. Their main argument is that Hawking’s result completely disregards the backreaction of the emission on the black hole itself. Refs. [19, 18] then suggest that this effect should no longer be negligible when the energy of the emitted quanta becomes comparable to that of the black hole itself. The proposed backreaction leads to a “memory burden” [19], which effectively suppresses the evaporation rate by inverse powers of the black hole entropy $S^{-n}$. This dependence of the suppression on the black hole entropy was conjectured in Ref. [18] and is motivated by the consideration that a black hole, viewed as a quantum system with maximal memory storage capacity, carries a high load of quantum information. When the black hole emits a quantum of Hawking radiation, the information stored previously in a set of gapless “memory” modes has to be re-written to the remaining internal degrees of freedom, which over time becomes increasingly costly in energy as fewer such gapless modes are available. Thus, naively one expects that the suppression of the evaporation rate should be related to the amount of information stored within the black hole, which is exactly what the Bekenstein-Hawking entropy $S$ measures. The free parameter $n>0$ quantifies the precise strength of the suppression. Since for $M>M_{\rm pl}$ the entropy of a black hole is “large”, the memory burden can significantly extend the black hole’s lifetime. Interestingly, Refs. [20, 21, 22] recently noted that, in the presence of the memory burden, PBHs with masses below $10^{10}\rm g$ could still account for the whole dark matter.
@@
In this work, we focus on even lighter PBHs which may have dominated the energy density of the universe in its very early stages and subsequently led to reheating via the emission of thermal radiation through Hawking evaporation [23, 24, 25, 26, 27, 28, 29]. Crucially, as shown in Refs. [30, 31, 32, 33, 34, 35, 36, 37, 38] (see also Refs. [39, 40] for reviews), a PBH dominated stage in the very early universe could lead to an observable GW background. This signature would result from perturbations induced by PBH shot noise at formation, while second-order induced GWs associated with the PBH formation mechanism would fall at frequencies above the LIGO/Virgo/Kagra band [41] in the so-called ultra-high frequency range [42, 43]. Since the GW production in this scenario depends on the dynamics before and after PBH evaporation, it may represent a unique opportunity to test modifications of Hawking evaporation. In this work, we will show that GWs induced right after evaporation by PBH number density fluctuations have the potential to probe the memory burden model. Also, since the main effect of the memory burden is to extend the duration of the PBH domination without drastically altering the final instants of evaporation, the GW signal is enhanced with respect to the standard case.
@@ II Memory burden and the PBH dominated universe
where $M_{\rm mb}=qM_{\rm f}$ parametrizes the moment when the modification to the semi-classical approximation sets in, with $M_{\rm f}$ being the PBH mass at formation and $q$ a free parameter. The semi-classical (sc) evaporation is recovered by taking the limit $n\to 0$. Note that from now on we use the subscript “f” to denote evaluation at the time of formation. We introduced ${\cal M}$ to distinguish two different possibilities discussed in Ref. [22]:
@@
where $g_{H}(T_{\rm PBH})$ are the spin-weighted degrees of freedom heavier than the PBH temperature $T_{\rm PBH}=M_{\rm pl}^{2}/M$. In what follows, we only consider ultra-light PBHs with $M\ll 10^{11}\,{\rm g}$ and, therefore, we use the high temperature SM relativistic degrees of freedom $g_{H}(T_{\rm PBH})\approx 108$, unless stated otherwise. It should be noted at this point that the black hole entropy (2) is generally determined by the surface area $A_{h}$ of the event horizon as ${S=A_{h}/(4G_{N})}$. In the case of a rotating and/or charged black hole, $A_{h}$ and, therefore, also $S$ depend on the charge $Q$ and the angular momentum $J$ in addition to the mass $M$. For simplicity, we will focus on non-rotating and non-charged PBHs in this work.
@@
We note that our results are also applicable to the scenarios considered in Ref. [12]. In particular, we can underline a few relevant examples. We introduce $\varepsilon$ which parametrizes, in their notation, the departure from the standard Hawking evaporation of a Schwarzschild black hole leading to a temperature $T\sim\frac{[\varepsilon/(1+\varepsilon)^{2}]}{M_{\rm PBH}}$. In the case of a nearly extremal Reissner-Nordstrom charged black hole, $\varepsilon=\sqrt{1-Q^{2}/M^{2}}\ll 1$, where $Q$ represents the PBH charge. To consider these scenarios, it is sufficient to assume $q=1$ while identifying the factor $\alpha^{-1}q^{3+2n}S_{\rm f}^{n}$ with $\frac{(1+\varepsilon)^{6}}{2^{6}\varepsilon^{4}}$. Another notable example is provided by higher dimensional black holes [14, 15]. An initial BH of size $r_{s}$ much larger than the size of extra dimensions $R$ evaporates following the 4D decay equation until its horizon becomes $r_{s}\simeq R$. Afterward, it would continue its decay at a rate that is slower than 4D due to the effect of the extra dimension. We can derive the corresponding phenomenological quantities by relating Eq. (4) to Eq. (A22) in [12]. As a final example, also a large number of additional degrees of freedom beyond the Standard Model could lead to a modified evaporation rate at small PBH masses, which we can describe within our setup by replacing Eq. (1) with Eq. (14) from [46].
@@ II.1 Extended PBH reheating
Let us now turn to the PBH reheating scenario (see e.g. Refs. [30, 32, 33] for more details on the standard case). We assume that PBHs form from the collapse of large primordial fluctuations after inflation in the radiation dominated universe [5, 6, 7, 8, 9, 10]. [^1] Note that PBHs could also form from first order phase transitions [47, 48, 49], the collapse of Q-balls [50, 51, 52], fifth forces [53, 54, 55] and preheating [56]. For analytical simplicity, we further assume a monochromatic PBH mass spectrum (for a broad mass spectrum in the PBH reheating scenario, see Ref. [35]) with the PBH mass given by
@@
$\Gamma=-\frac{d\log M}{dt}=\begin{cases}\frac{1}{3(t_{\rm eva}^{\rm sc}-t)}\qquad{\rm for}\quad t\leq t_{\rm mb}\\ \frac{1}{3\alpha(t_{\rm eva}-t)}\qquad{\rm for}\quad t>t_{\rm mb}\end{cases}~{}.$ (12)
@@
In general, we find two possibilities depending on whether the early PBH-radiation equality (which we denote as “eeq” and which is defined by $\rho_{\rm PBH}(t_{\rm eeq})=\rho_{\rm rad}(t_{\rm eeq})$) happens before or after the memory burden time $t_{\rm mb}$ (7). In the first case, that is $t_{\rm mb}<t_{\rm eeq}$, the initial fraction of the PBHs is too low to reach matter radiation equality before the memory burden time. In fact, this case roughly corresponds to the standard scenario where PBHs evaporate before dominating the universe. Now, due to the memory burden, PBHs have a second chance to dominate the universe. In the second case, where $t_{\rm mb}>t_{\rm eeq}$, the early PBH-radiation equality occurs before the memory burden time. Then, as seen in Fig. 1, there is a short period in the PBH dominated stage when the PBH energy density decays, but PBHs still dominate the universe. We find that, in that case, $\Omega_{\rm PBH}(t_{\rm mb})\approx 0.59$ for $q=1/2$. It should be noted that the precise value of $\Omega_{\rm PBH}(t_{\rm mb})$ depends on $q$. For example, for $q\lesssim 0.41$ we find that $\Omega_{\rm PBH}(t_{\rm mb})<1/2$ and radiation temporarily dominates the universe. Depending on the parameters of the modified evaporation, we find that for small enough values of $q$ PBHs do not come to dominate again after $t_{\rm mb}$, while for other choices there is a second stage of PBH domination. We do not consider these possibilities here and focus on the regime $10.41<q<$ 1. In particular, we will take $q=\frac{1}{2}$ for all plots, unless stated otherwise. Also, note that the PBH remnant scenario would correspond to $q\sim M_{\rm pl}/M_{\rm f}$ (and $n\to\infty$ so that Hawking evaporation stops completely).
@@
$M_{\rm f}<M^{\rm mb}_{\rm f,max}=5\times 10^{8}\,{\rm g}\left(\frac{e^{12n}q^{n+\frac{3}{2}}7^{\frac{2n}{3}}10^{8n}}{\alpha}\right)^{-\frac{2}{2n+3}}\approx 5\times 10^{8}\,{\rm g}\left(\frac{e^{31.7n}q^{n+\frac{3}{2}}}{\alpha}\right)^{-\frac{2}{2n+3}}~{}.$ (19)
@@
Lastly, we can express the BBN bound in terms of an upper bound on $n$ by enforcing the Hubble parameter at formation to be smaller than the aforementioned bound $H_{\rm f}<2.5\times 10^{-5}\,M_{\rm pl}$ [58], which provides a lower bound for the mass which is roughly given by $M_{\rm f,max}^{\rm mb}\gtrsim 1\,{\rm g}$. Thus, we find
$n<\frac{59-3\log q}{2(12+\log q)}$ (20)
@@ III Memory burden and the induced GW signal from PBH density fluctuations
We proceed to estimate the GW spectrum induced by PBH number density fluctuations. We will closely follow Refs. [33, 39], which builds upon the work of Refs. [30, 63, 64, 31, 32], adapting the relevant calculations to the current case. The main effect of the modification to Hawking evaporation given by Eqs. (4) is an extension of the PBH dominated era, proportional to the PBH entropy. Thus, one naively expects that the amplitude of the GW spectrum is enhanced and that the peak moves to lower frequencies with respect to the standard case.
@@
$\frac{k_{\rm uv}}{k_{\rm eva}}\approx 4989\,\alpha^{-1/3}e^{8n}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2}{3}(1+n)}\,.$ (22)
@@
${f_{\rm uv}}\approx 3.7\times 10^{6}\,{\rm Hz}\,\alpha^{1/6}e^{-4n}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{-\frac{5}{6}-\frac{n}{3}}\left(\frac{g_{*}(T_{\rm eva})}{106.75}\right)^{1/4}\left(\frac{g_{*s}(T_{\rm eva})}{106.75}\right)^{-1/3}\,.$ (24)
@@
${\cal P}_{\Phi}(k)\Big{|}_{\rm eva}={\cal A}_{\Phi}\left(\frac{k}{k_{\rm uv}}\right)^{m}\Theta(k_{\rm uv}-k)\,,$ (25)
@@
We will use $\Omega_{\rm GW,\rm res}^{\rm peak}$ from Eq. (III) as a good approximation for the peak amplitude of the GW spectrum. It should be noted that in the case when $\beta<\beta_{*}$, there is an additional suppression factor $q^{8}$ coming from $k_{\rm eeq}$ in Eq. (III). Taking $q\sim 1/2$ results in a suppression of $0.004$, which would be exactly compensated by taking a factor $3$ larger value for $\beta$. The GW spectrum evaluated today is then given by
$\Omega_{\rm GW,0}h^{2}$ $\approx 1.62\times 10^{-5}\left(\frac{\Omega_{r,0}h^{2}}{4.18\times 10^{-5}}\right)\left(\frac{g_{*}(T_{\rm eva})}{106.75}\right)\left(\frac{g_{*s}(T_{\rm eva})}{106.75}\right)^{-4/3}\Omega_{\rm GW,eva}(f)\,.$ (32)
@@
Figure 5: The induced gravitational wave spectrum (32) for PBH mass ${M_{\rm f}=2\times 10^{3}$\mathrm{g}$}$, $\beta=10^{-8}$ and $n=\frac{1}{2}$ in the scenario mb1 ($\alpha=1$) (solid red line), as well as the spectrum for scenario mb2 ($\alpha=1+2n/3$) with a PBH mass ${M_{\rm f}=6\times 10^{3}$\mathrm{g}$}$, $\beta=5\times 10^{-13}$ and $n=1$ (dashed red line). As one can see, for this choice of parameters, the signals enter the observational windows for several of the next-generation GW detectors. We also show the power-law integrated sensitivity curves as in Ref. [67] (but also see Ref. [68]) for LIGO A+, Voyager, Einstein Telescope (ET) (triangular configuration with 15km arms [69]), Cosmic Explorer (CE) and DECIGO experiments (for the sensitivity curves see Refs. [70, 71, 72, 68, 73, 74]). In purple, we show the upper bounds from the search for an isotropic GW background from the LVK collaboration [75]. The current (integrated) constraint from BBN [76, 77, 78] is shown with a dashed blue line.
@@
$\Omega_{\rm GW,\rm res}^{\rm peak}\approx 1.2\times 10^{15}\,\beta^{16/3}q^{4}\left(\frac{\sqrt{2}\alpha}{\sqrt{3\alpha-1}}\right)^{-\frac{4}{3\alpha}}\frac{e^{\frac{8}{\alpha}\left[{(7\alpha-\frac{4}{3})n+\frac{17}{12}(\alpha-1)}\right]}}{\alpha^{\frac{7}{3}-\frac{4}{9\alpha}}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2(21\alpha-4)(1+n)}{9\alpha}}\,.$ (34)
@@
$\beta<1.2\times 10^{-4}\,q^{-3/4}\frac{\alpha^{\frac{1}{6\alpha}+\frac{7}{16}}}{(3\alpha-1)^{\frac{1}{8\alpha}}}e^{\frac{-84\alpha n+16n+17}{8\alpha}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{-\frac{(21\alpha-4)(1+n)}{24\alpha}}\,.$ (35)